T= -0.2x^2+12x+11
the questions are:
what is the point of the vertex? what do the soulations represent?

Answer by stanbon(75887)

T= -0.2x^2+12x+11
the questions are:
what is the point of the vertex? what do the solutions represent?
--------------
Vertex occurs where x = -b/2a = -12/(2*-0.2) = -12/-0.4 = 30
Then T(30) = -0.2(30)^2 + 12(30) + 11 = 191
Vertex(30,191)
---------------------
Solutions are values of "x" that satisfy the equation
-0.2x^2+12x+11 = 0
They are the point(s) where the graph of the equation
crosses the x-axis.
=========================
Cheers,
Stan H.